OpenStudy (anonymous):

PLEASE HELP A.S.A.P! Rectangle A has an area of 4 - x^2. Rectangle B has an area of x^2 + 2x - 8. In simplest form, what is the ratio of the area of Rectangle A to the area of Rectangle B? Please show your work so that I can understand:)

OpenStudy (kropot72):

The numerator and the denominator both can be factorised: \[\frac{(2+x)(2-x)}{(x+4)(x-2)}\] Multiply numerator and denominator by -1 to enable cancellation: \[\frac{(2+x)(x-2)}{(x+4)(x-2)}\] Answer is: \[-\frac{2+x}{4+x}\]

OpenStudy (anonymous):

Thank you so much for your help!!!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

Could you explain to me real quick how you changed it from: 2+x)(2−x) over (x+4)(x−2) to (2+x)(x−2) over (x+4)(x−2)?? I'm kinda confused:/

OpenStudy (anonymous):

Never mind I mean the factoring part! I dont understand facotring very well

OpenStudy (kropot72):

The numerator of the original expression is the difference of two squares.To facorise you take the square root of each term and arrange as follows: 4 - x^2 = (2 + x)(2 - x) To factorise the denominator you need to find factors of 8 that have a difference of 2. These are 4 and 2. x^2 + 2x - 8 = (x + 4)(x - 2)